OpenStudy (mackenzie2013):

A moving company charges a $2,000 fee plus $100 per hour. Another moving company charges an $1,800 fee plus $150 per hour. How long is a job for which the two companies' costs are the same? 2 hours 5 hours 3 hours 4 hours

OpenStudy (ibrafacts):

u sure this si written right?

OpenStudy (amr979):

after 4 hours company 1 = 2000+100*4=2400 and company 2 = 1800+150*4=2400